solve D.E. y''=1+(y' )^2
1
Expert's answer
2020-03-20T17:22:29-0400

"y'' = 1 + (y')^2\\
\frac{dy'}{dx} = 1 + (y')^2\\
\frac{dy'}{1 + (y')^2} = dx\\
x = \intop\frac{dy'}{1 + (y')^2} = arctan(y') + c_1\\
y' = tan(x - c_1)\\
y = \intop tan(x - c_1)dx = \intop \frac{sin(x - c_1)}{cos(x - c_1)}d(x - c_1) = \intop - \frac{dcos(x - c_1)}{cos(x - c_1)} = - ln|cos(x - c_1)| + c_2"